a car parked near th edge falls from a hill and hits the ground in 0.5s.
1.what is the speed of car when it hits ground?
2.what is its average speed during its fall?
3.how high is the hill from the ground?

Grade:9

1 Answers

Arun
6 years ago
Dear student
 

Given: u = 0m/s, a = 10m/s2, t = 0.5s

1) v = u+at

=> 0+ (10)(0.5)

=> 5m/s

Therefore, final velocity = 5m/s

2) vavg = (u+v)/ 2

=> (0+ 5)/2

=> 5/2

=>  Average velocity is 2.5m/s

3) v2 - u2 = 2as

=> s = (v- u2)/2a

=>5- 02/ 2(10)

=> (25 - 0)/20

=>25/20

=>  height of ledge is 1.25m